A chemist needs 2.38 grams of a liquid compound with a density of 0.718 g/cm3. What volume of the compound is required?
1
Expert's answer
2019-07-03T05:55:34-0400

V = m/d = 2.38g/(0.718g/cm3) = 3.315 cm3